OpenStudy (anonymous):

|3(6 - 15) + |3 - 22|| it needs to be simplified

OpenStudy (mathmale):

Mind taking a moment to explain what (6-15) means? and then explaining what |3-22| means? Once I can verify that you understand this material, we can move forward.

OpenStudy (anonymous):

im confused how do u want me to explain it?? mathmale

OpenStudy (mathmale):

@jaxsonH22: thanks for your contribution, but our job is NOT, repeat, NOT, to give out free answers. If you can lead someone else towards finding his or her own solution, please do so; otherwise, please stay silent.

OpenStudy (mathmale):

"Order of Operation" rules state that we must do any work inside parentheses FIRST, before anything else. If you're given -(6-15), you must evaluate 6-15 first, because it's inside parentheses. What is 6-15?

OpenStudy (anonymous):

uhhhhh..........thanks @JaxsonH22 and @mathmale it is -9

OpenStudy (mathmale):

@jaxsonh22: Rest assured that you haven't made an enemy! It's just that so many people join OpenStudy without reading the Code of Conduct first, and don't realize that the code forbids either asking for or providing just answers. I do hope to work with you constructively in the future.

OpenStudy (mathmale):

that's right, it's -9. Therefore, -(6-15) is -(-9) = ?

OpenStudy (anonymous):

Same here @mathmale

OpenStudy (anonymous):

-9^5

OpenStudy (mathmale):

-(6-15) can be explained in words as follows: "First evaluate the expression 6-15 within the parentheses. That's -9. Next, take the negative of that (-9). Result?

OpenStudy (anonymous):

uhhh..............positive 9

OpenStudy (mathmale):

Perfect! Now let's move on to -|3-22|. That symbol | | means "absolute value." Are you familiar with "absolute value?"

OpenStudy (mathmale):

If so, what is the absolute value of (3-22)?

OpenStudy (anonymous):

19? cause i have to solve in the perentces first

OpenStudy (mathmale):

Yes, you have to do the work inside the ( ) first. You've done that correctly, obtaining -19, and then you took the "absolute value" of that, |-19| = 19. Very good!

OpenStudy (anonymous):

thanks

OpenStudy (mathmale):

Going back to the original problem: + |3 - 22| Please evaluate both of these: + |3 - 22| - |3 - 22| Hint: Think carefully ab out what "absolute value" means, and remember Order of Operations rules.

OpenStudy (mathmale):

Hint: You have already correctly finished some of the work required here.

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

0

OpenStudy (mathmale):

Actually, I meant to ask you to evaluate two separate expressions. + |3 - 22| = + (19) = ?? - |3 - 22| = - ( ) = ?

OpenStudy (mathmale):

This is just practice; we'll solve your original problem next!

OpenStudy (anonymous):

kk uhhh...........-24

OpenStudy (mathmale):

+ |3 - 22| = + (19) = 19. Try again on: - |3 - 22| = - ( ) = ?

OpenStudy (anonymous):

positive 24

OpenStudy (mathmale):

- |3 - 22| = - ( 19 ) = ?? Note that |3-22| is +19

OpenStudy (anonymous):

-25

OpenStudy (anonymous):

-25= 19 and 19= 25

OpenStudy (mathmale):

Just take the negative of 19. It's -19.

OpenStudy (mathmale):

I'm sure you recognize that -25 is not equal to 19. Let's go back to the original question: |3(6 - 15) + |3 - 22|| it needs to be simplified

OpenStudy (anonymous):

forget it im not smart enough byee @mathmale

OpenStudy (mathmale):

Notice how there are absolute value signs around the whole expression 3(6 - 15) + |3 - 22|

OpenStudy (mathmale):

@obbey_jaz10934 I'm positive you can do this problem and understand it as well. Please stick with me. If you can't do that now, come back later when you're feeling up to it.

OpenStudy (anonymous):

k

OpenStudy (mathmale):

Now or later?

OpenStudy (mathmale):

3(6 - 15) + |3 - 22| = 3(-9) + |-19} -27 + ? Please answer this when you're ready, @obbey_jaz10934